# Regularity Propagation of Global Weak Solutions to a Navier–Stokes–Cahn–Hilliard System for Incompressible Two-phase Flows with Chemotaxis and Active Transport

### v1

@@ Abstract
We analyze a diffuse interface model that describes the dynamics of incompressible viscous two-phase flows, incorporating mechanisms such as chemotaxis, active transport, and long-range interactions of Oono’s type. The evolution system couples the Navier–Stokes equations for the volume-averaged fluid velocity $\bm{v}$, a convective Cahn–Hilliard equation for the phase-field variable $\varphi$, and an advection-diffusion equation for the density of a chemical substance $\sigma$. For the initial-boundary value problem with a physically relevant singular potential in three dimensions, we demonstrate that every global weak solution $(\bm{v},\varphi,\sigma)$ exhibits a propagation of regularity over time. Specifically, after an arbitrary positive time, the phase-field $\varphi$ transitions into a strong solution, whereas the chemical density $\sigma$ only partially regularizes. Subsequently, the velocity field $\bm{v}$ becomes regular after a sufficiently large time, followed by a further regularization of the chemical density $\sigma$, which in turn enhances the spatial regularity of $\varphi$. Furthermore, we show that every global weak solution stabilizes towards a single equilibrium as $t\to+\infty$. Our analysis uncovers the influence of chemotaxis, active transport, and long-range interactions on the propagation of regularity at different stages of time. The proof relies on several key points, including a novel regularity result for a convective Cahn–Hilliard–diffusion system with minimal regularity requirements on the velocity field $\bm{v}$, the strict separation property of $\varphi$ for large times, as well as two conditional uniqueness results pertaining to the full system and its subsystem for $(\varphi,\sigma)$ with a given velocity, respectively.
@@ 1 Introduction
The study of multi-phase flows has garnered significant attention in various fields of science and engineering. One of the efficient mathematical approaches is the diffuse interface model, which assumes that an immiscible fluid mixture undergoes a smooth and rapid transition within an interfacial region between its components, known as partial diffusive mixing. This approach results in an Eulerian formulation of the phase-field equation and permits topology changes such as breakup and coalescence. Different types of diffuse interface models have been introduced to accommodate a wide range of physical and biological applications, see e.g., [3, 5, 13, 26, 19, 31, 37, 39] and the references cited therein.
@@
System (1.1) can be regarded as a simplified version of the general thermodynamically consistent diffuse interface model derived in [36]. This model describes the dynamics of a mixture of two viscous incompressible fluids, with a chemical species subject to the diffusion process, along with significant mechanisms such as chemotaxis and active mass transport (see also [43] and the references cited therein). The parameter $\varepsilon>0$ characterizes the thickness of interfacial layers between two fluid components. Since the asymptotic behavior as $\varepsilon\to 0^{+}$ (i.e., the sharp-interface limit) will not be investigated in this study, we simply set $\varepsilon=1$. The state variables of system (1.1) are given by $(\bm{v},p,\varphi,\mu,\sigma)$. The vectorial function $\bm{v}:\Omega\times[0,+\infty)\to\mathbb{R}^{3}$ denotes the volume-averaged velocity with $D\bm{v}=\frac{1}{2}(\nabla\bm{v}+(\nabla\bm{v})^{\mathrm{T}})$ being the symmetrized velocity gradient, and the scalar function $p:\Omega\times[0,+\infty)\to\mathbb{R}$ denotes the (modified) pressure. The order parameter (also called the phase-field) $\varphi:\Omega\times[0,+\infty)\to[-1,1]$ denotes the difference in volume fractions of the binary mixture such that the region $\{\varphi=1\}$ represents the fluid 1 and $\{\varphi=-1\}$ represents the fluid 2 (i.e., the values ±$\pm 1$ represent the pure phases). The scalar function $\sigma:\Omega\times[0,+\infty)\to\mathbb{R}$ denotes the concentration of an unspecified chemical species (e.g., a nutrient in the context of tumor growth modelling [19, 18]). Finally, the scalar function $\mu:\Omega\times(0,+\infty)\to\mathbb{R}$ is referred to as the chemical potential in the terminology of Cahn–Hilliard [9]. For the sake of simplicity, we assume that the density difference between the two components of the fluid mixture is negligible and set the density to be one. Additionally, we assume that the mobilities for $\varphi$ and $\sigma$ are positive constants (set to be one), while allowing the fluid mixture to have unmatched viscosities. Assuming that $\nu_{1}$, $\nu_{2}>0$ are viscosities of the two homogeneous fluids, the mean viscosity is given by the concentration dependent function $\nu=\nu(\varphi)$, for instance, the linear combination:
@@
or by some more general polynomial functions (see [39]). This approximation rules out possible singularities of the logarithmic potential (1.5) (and its derivatives) at ±$\pm 1$ and brings great convenience in the mathematical analysis as well as numerical simulations for the Cahn–Hilliard type equations, see [6, 12, 15, 33, 37, 46] and the references therein.
The evolution of velocity $\bm{v}$ and pressure $p$ is governed by the Navier–Stokes equations under the influence of a Korteweg-type force proportional to $(\mu+\chi\sigma)\nabla\varphi$. The dynamics of $(\varphi,\mu)$ are governed by a convective Cahn–Hilliard equation, while $\sigma$ satisfies an advection-diffusion equation. In (1.1c) and (1.1d) we denote by $\overline{\varphi}=|\Omega|^{-1}\int_{\Omega}\varphi(x)\,\mathrm{d}x$ the mean of the phase function $\varphi$ over $\Omega$. Besides, $\mathcal{N}$ denotes the inverse of the minus Laplacian operator subject to the homogeneous Neumann boundary condition on some function space with zero mean constraint (see Section 2). As shown in [36], the hydrodynamic system (1.1)–(1.3) is thermodynamically consistent. Sufficiently regular solutions satisfy the energy dissipation law
@@
Coupling structures of the system (1.1) are characterized by the parameters $\alpha$, $\beta$ and $\chi$. The two terms $\alpha(\overline{\varphi}-c_{0})$ and $\beta\mathcal{N}(\varphi-\overline{\varphi})$ represent certain nonlocal interaction (i.e., reaction) between the two fluid components. Throughout the paper, we assume that $\alpha\geq 0$ and $\beta\in\mathbb{R}$. The role of $\alpha(\overline{\varphi}-c_{0})$ has two aspects. In view of (1.7), it can influence the energy dissipation of the system. On the other hand, it may yield different mass relations for $\varphi$, that is, if $\alpha=0$ or $\overline{\varphi_{0}}=c_{0}$ for $\alpha>0$, then the mass $\overline{\varphi}(t)$ is conserved in time, otherwise, $\overline{\varphi}(t)$ converges exponentially fast to $c_{0}$ provided that $\alpha>0$ (i.e., the so-called off-critical case, cf. [6, 22, 40]). Neglecting the coupling with $\bm{v}$ and $\sigma$, for $\alpha=\beta=0$, we recover the classical Cahn–Hilliard equation with constant mobility [9, 10, 39]. When taking $\alpha=\beta>0$, we obtain the well-known Cahn–Hilliard–Oono (CHO) equation
@@
$\mathcal{F}_{\mathrm{OK}}(\varphi)=\int_{\Omega}\left(\frac{1}{2}|\nabla\varphi|^{2}+\varPsi(\varphi)\right)\mathrm{d}x+\alpha\int_{\Omega}\!\int_{\Omega}G(x-y)(\varphi(x)-\overline{\varphi})(\varphi(y)-\overline{\varphi})\,\mathrm{d}x\mathrm{d}y,$
@@
where $G$ denotes the Green function associated to the minus Laplacian subject to a homogeneous Neumann boundary condition. For results on well-posedness and long-time behavior of solutions to the CHO equation, we refer to, e.g.,[22, 38].
@@
The coefficient $\chi$ is related to some specific transport mechanisms such as chemotaxis (as seen in (1.1d)) and active transport (as seen in (1.1e)). In the context of tumor growth modelling, $\chi$ models the chemotactic response of the mixture towards the nutrient, as well as the evolution of the nutrient influenced by the mixture. For detailed explanations, see [19, 36]. The term “active transport" is used when $\chi>0$ in the biological sense, indicating that some kind of mechanism is required to maintain the transport. This is in contrast to passive transport processes, such as diffusion, which are driven only by the concentration gradient [18]. From the mathematical perspective, the term $-\chi\Delta\varphi$ introduces a cross diffusion structure in (1.1e), which poses additional challenges in the analysis. However, the sign of $\chi$ does not play a role, so we allow $\chi$ to be any real number, i.e., $\chi\in\mathbb{R}$.
@@
We observe that the well-known “Model H” for incompressible viscous two-phase flows (see [31, 26]) can be recovered from the system (1.1) by neglecting $\sigma$ and setting $\alpha=\beta=\chi=0$. According to (1.7) and (1.9), the resulting Navier–Stokes–Cahn–Hilliard (NSCH) system possesses a dissipative structure (i.e., its total energy decreases over time) and conserves the $\overline{\varphi}$ for all $t\geq 0$. Due to these two crucial properties, the NSCH system (with either a regular or a singular potential) has been extensively studied in the literature. In this regard, we refer to [1, 7, 15, 23, 37, 46] and the references therein for results on well-posedness, long-time behavior of global solutions, and asymptotic analyses. See also [3, 2, 20, 21] for contributions on the more complex case with unmatched densities. When the equation for $\sigma$ is neglected, our system (1.1) also reduces to a variant of the Navier–Stokes–Cahn–Hilliard–Oono (NSCHO) system, which has been studied in [6, 40, 32]. Compared to the basic energy law (1.7), the NSCHO system in [6, 40] exhibits a higher-order energy production rate without a definite sign, given by $-\alpha\int_{\Omega}(\varphi-c_{0})\widetilde{\mu}\,\mathrm{d}x$ with $\widetilde{\mu}=-\Delta\varphi+\varPsi^{\prime}(\varphi)$ (recalling that $\overline{\varphi}$ depends solely on time and decays to $c_{0}$ exponentially fast). This fact leads to significant challenges in studying the global well-posedness and long-time behavior, as discussed in [40]. For the single CHO equation (1.11) without fluid interaction, the authors of [22] overcame the difficulty by considering the Ohta–Kawasaki functional $\mathcal{F}_{\mathrm{OK}}$ and rewriting the CHO equation as
@@
Thanks to the energy balance (1.7) and mass relations (1.9), (1.10), in our recent work [30], we were able to show that, in the three-dimensional setting, if the initial velocity is small and both the initial phase-field function and the initial chemical density are small perturbations of a local minimizer of the free energy $\mathcal{F}$, then problem (1.1)–(1.3) admits a unique global strong solution. Furthermore, we proved the uniqueness of the asymptotic limit for every global strong solution as $t\to+\infty$ and provided an estimate on the convergence rate. The proofs relied on the dissipative structure of the system and an extended Łojasiewicz–Simon type gradient inequality. However, whether global weak solutions to problem (1.1)–(1.3) exist and will regularize like solutions to the NSCH system, as demonstrated in [1, 2], remains an open question.
@@
(1) Our first result states that every global weak solution $(\bm{v},\varphi,\sigma)$ exhibits a propagation of regularity over time, see Theorem 2.1. Specifically, after an arbitrary positive time, the phase-field $\varphi$ transitions into a strong solution, whereas the chemical density $\sigma$ only partially regularizes. Subsequently, the velocity field $\bm{v}$ becomes regular after a sufficiently large time, followed by a further regularization of the chemical density $\sigma$, which in turn enhances the spatial regularity of $\varphi$. The proof of Theorem 2.1 is inspired by the recent work [2] on the global regularity of weak solutions to the NSCH system with unmatched densities. While we tackle the simpler case of a fluid mixture with constant density, additional challenges arise due to chemotaxis, mass transport and nonlocal interactions, making it difficult to establish the regularity of $(\varphi,\sigma)$. Specifically, the chemical density $\sigma$ exhibits a novel phenomenon, undergoing regularization at different stages of time. For discussions on the strategy of proofs, we refer to Section 2.3.
@@
(2) Our second result pertains to the long-time behavior. Under the additional assumption that $\varPsi$ is real analytic in $(-1,1)$ (which holds true for (1.5)), we demonstrate that every global weak solution $(\bm{v},\varphi,\sigma)$ converges to a single equilibrium $(\mathbf{0},\varphi_{\infty},\sigma_{\infty})$ as $t\to+\infty$ and provide an estimate of the convergence rate, as stated in Theorem 2.2. This conclusion is reached by applying Theorem 2.1 and employing the same argumentation as presented in [30].
@@
Before concluding the Introduction, we review some related results in the literature. With a regular potential, including (1.6), and some more general reaction terms, problem (1.1)–(1.3) was first studied in [36]. In this work, the authors proved the existence of global weak solutions in both two and three dimensions, and the existence of a unique global strong solution in two dimensions. However, due to the coupling between $\varphi$ and $\sigma$, as well as the loss of maximum principle for the Cahn–Hilliard equation with regular potentials, they imposed certain technical assumption on the coefficients to achieve their goals. Later, in [27], this restriction was removed by considering a singular potential like (1.5), which is more physically relevant. Using a semi-Galerkin approach, the author proved the existence of global weak solutions in both two and three dimensions, as well as the uniqueness of global weak solutions in two dimensions. For the same system, the authors of [29] further established the global strong well-posedness for arbitrarily large regular initial data in two dimensions. The arguments in [27, 29] rely on the key property that the singular potential $\varPsi$ ensures that the phase function $\varphi$ remains within the physical interval $[-1,1]$ along the time evolution. Moreover, in the two-dimensional case, $\varphi$ is strictly separated from the pure states ±$\pm 1$ for $t>0$. This separation property plays a crucial role in studying the regularity and long-time behavior of solutions to the Cahn–Hilliard equation with a singular potential, see [4, 10, 22, 39, 41]. We also refer to [1, 7, 11, 20, 24, 23] for generalized systems with fluid coupling. Finally, we note that Oono’s linear term represents possibly the simplest case of (nonlocal) reaction. Our results provide answers to some open questions raised in [40] concerning the global regularity and asymptotic stabilization of global weak solutions to the NSCHO system with singular potential. On the other hand, more general reaction terms have been integrated into the Cahn–Hilliard equation for biological applications [14, 35], as well as into extended systems for tumor modelling [17, 18, 19, 34]. Investigating the global well-posedness and long-time behavior of global weak/strong solutions to the hydrodynamic system (1.1)–(1.3) with these general reaction terms constitutes an interesting and challenging task.
@@
Plan of the paper. The remainder of this paper is organized as follows. In Section 2, we introduce the mathematical settings and state the main results. In Section 3 we analyze a convective Cahn–Hilliard–diffusion system with a prescribed divergence-free velocity, which may have its independent interest. Section 4 is devoted to the proofs of the main results.

### v2

@@ Abstract
We analyze a diffuse interface model that describes the dynamics of incompressible viscous two-phase flows, incorporating mechanisms such as chemotaxis, active transport, and long-range interactions of Oono’s type. The evolution system couples the Navier–Stokes equations for the volume-averaged fluid velocity $\bm{v}$, a convective Cahn–Hilliard equation for the phase-field variable $\varphi$, and an advection-diffusion equation for the density of a chemical substance $\sigma$. For the initial boundary value problem with a physically relevant singular potential in three dimensions, we demonstrate that every global weak solution $(\bm{v},\varphi,\sigma)$ exhibits a propagation of regularity over time. Specifically, after an arbitrary positive time, the phase-field variable $\varphi$ transitions into a strong solution, whereas the chemical density $\sigma$ only partially regularizes. Subsequently, the velocity field $\bm{v}$ becomes regular after a sufficiently large time, followed by a further regularization of the chemical density $\sigma$, which in turn enhances the spatial regularity of $\varphi$. Furthermore, we show that every global weak solution stabilizes towards a single equilibrium as $t\to+\infty$. Our analysis uncovers the influence of chemotaxis, active transport, and long-range interactions on the propagation of regularity at different stages of time. The proof relies on several key points, including a novel regularity result for a convective Cahn–Hilliard–diffusion system with a velocity field $\bm{v}$ of Leray type, the strict separation property of $\varphi$ for large times, as well as two conditional uniqueness results pertaining to the full system and its subsystem for $(\varphi,\sigma)$ with a given velocity, respectively.
@@ 1 Introduction
The study of multi-phase flows has garnered significant attention in various fields of science and engineering. One of the efficient mathematical approaches is the diffuse interface model, which assumes that an immiscible fluid mixture undergoes a smooth and rapid transition within an interfacial region between its components, known as partial diffusive mixing. This approach results in an Eulerian formulation of the phase-field equation and permits topology changes such as breakup and coalescence. Different types of diffuse interface models have been introduced to accommodate a wide range of physical and biological applications, see, e.g., [4, 7, 16, 30, 22, 36, 43, 45] and the references cited therein.
@@
System (1.1) can be regarded as a simplified version of the general thermodynamically consistent diffuse interface model derived in [41]. This model describes the dynamics of a mixture of two viscous incompressible fluids, with a (soluble) chemical species subject to the diffusion process, along with significant mechanisms such as chemotaxis and active mass transport (see also [50] and the references cited therein). The parameter $\varepsilon>0$ characterizes the thickness of interfacial layers between two fluid components. Since the asymptotic behavior as $\varepsilon\to 0^{+}$ (i.e., the sharp-interface limit) will not be investigated in this study, we simply set $\varepsilon=1$ in the subsequent analysis. The state variables of system (1.1) are given by $(\bm{v},p,\varphi,\mu,\sigma)$. The vectorial function $\bm{v}:\Omega\times[0,+\infty)\to\mathbb{R}^{3}$ denotes the volume-averaged velocity with $D\bm{v}=\frac{1}{2}(\nabla\bm{v}+(\nabla\bm{v})^{\mathrm{T}})$ being the symmetrized velocity gradient, and the scalar function $p:\Omega\times[0,+\infty)\to\mathbb{R}$ denotes the (modified) pressure. The order parameter (also called the phase-field variable) $\varphi:\Omega\times[0,+\infty)\to[-1,1]$ denotes the difference in volume fractions of the binary mixture such that the region $\{\varphi=1\}$ represents the fluid 1 and $\{\varphi=-1\}$ represents the fluid 2 (i.e., the values $\pm 1$ represent the pure phases). The scalar function $\sigma:\Omega\times[0,+\infty)\to\mathbb{R}$ denotes the concentration of an unspecified chemical species (e.g., a nutrient in the context of tumor growth modeling [22, 21]). Finally, the scalar function $\mu:\Omega\times(0,+\infty)\to\mathbb{R}$ is referred to as the chemical potential in the terminology of Cahn–Hilliard [11]. For the sake of simplicity, we assume that the density difference between the two components of the fluid mixture is negligible and set the fluid density to be one. Additionally, we assume that the mobilities for $\varphi$ and $\sigma$ are positive constants (set to be one), while allowing the fluid mixture to have unmatched viscosities. Assuming that $\nu_{1}$, $\nu_{2}>0$ are viscosities of the two homogeneous fluids, the mean viscosity is given by the concentration dependent function $\nu=\nu(\varphi)$, for instance, the linear combination:
@@
or by some more general polynomial functions (see [45]). This type of approximation rules out possible singularities of the logarithmic potential (1.5) (and its derivatives) at $\pm 1$ and brings great convenience in the mathematical analysis as well as numerical simulations for the Cahn–Hilliard equation and its variants, see [8, 15, 18, 38, 43, 54] and the references therein.
The evolution of velocity $\bm{v}$ and pressure $p$ is governed by the Navier–Stokes equations (1.1a)–(1.1b) under the influence of a Korteweg-type force proportional to $(\mu+\chi\sigma)\nabla\varphi$. The dynamics of $(\varphi,\mu)$ is governed by a convective Cahn–Hilliard equation (1.1c)–(1.1d), while $\sigma$ satisfies an advection-diffusion equation (1.1e). In (1.1c) and (1.1d) we denote by $\overline{\varphi}=|\Omega|^{-1}\int_{\Omega}\varphi(x)\,\mathrm{d}x$ the mean of the phase function $\varphi$ over $\Omega$. In addition, $\mathcal{N}$ denotes the inverse of the minus Laplacian operator subject to the homogeneous Neumann boundary condition in some function space with a zero mean constraint (see Section 2). As shown in [41], the hydrodynamic system (1.1)–(1.3) is thermodynamically consistent. Sufficiently regular solutions satisfy the law of energy dissipation
@@
The coupling structures of the system (1.1) are characterized by the parameters $\alpha$, $\beta$ and $\chi$. The two terms $\alpha(\overline{\varphi}-c_{0})$ and $\beta\mathcal{N}(\varphi-\overline{\varphi})$ represent certain nonlocal interactions (i.e., reactions) between the two fluid components. Throughout the paper, we assume that $\alpha\geq 0$ and $\beta\in\mathbb{R}$. The role of $\alpha(\overline{\varphi}-c_{0})$ has two aspects. In view of (1.7), it can influence the energy dissipation of the system. On the other hand, it may yield different mass relations for $\varphi$, that is, if $\alpha=0$ or $\overline{\varphi_{0}}=c_{0}$ for $\alpha>0$, then the mass $\overline{\varphi}(t)$ is conserved in time, otherwise $\overline{\varphi}(t)$ converges exponentially fast to $c_{0}$ provided that $\alpha>0$ (i.e., the so-called off-critical case, cf. [8, 26, 46]). Neglecting the coupling with $\bm{v}$ and $\sigma$, for $\alpha=\beta=0$, we recover the classical Cahn–Hilliard equation with constant mobility [11, 12, 45]. When taking $\alpha=\beta>0$, we obtain the well-known Cahn–Hilliard–Oono (CHO) equation
@@
$\mathcal{F}_{\mathrm{OK}}(\varphi)=\int_{\Omega}\left(\frac{1}{2}|\nabla\varphi|^{2}+\varPsi(\varphi)\right)\mathrm{d}x+\alpha\int_{\Omega}\!\int_{\Omega}G(x,y)(\varphi(x)-\overline{\varphi})(\varphi(y)-\overline{\varphi})\,\mathrm{d}x\mathrm{d}y,$
@@
where $G$ denotes the Green function associated to the (minus) Laplacian subject to a homogeneous Neumann boundary condition (see e.g., [35]). For results on well-posedness and long-time behavior of solutions to the CHO equation, we refer to, e.g., [26, 44].
@@
The coefficient $\chi$ is related to some specific transport mechanisms, such as chemotaxis (as seen in (1.1d)) and active transport (as seen in (1.1e)). In the context of tumor growth modeling, $\chi$ characterizes the chemotactic response of the mixture to the nutrient, as well as the evolution of the nutrient influenced by the mixture. For detailed explanations, see [22, 41]. The term “active transport" is used when $\chi>0$ in the biological sense, indicating that some kind of mechanism is required to maintain the transport. This is in contrast to passive transport processes, such as diffusion, which are driven only by the concentration gradient [21]. From a mathematical perspective, the term $\chi\Delta\varphi$ introduces a cross-diffusion structure in (1.1e), which poses additional challenges in the analysis. However, the sign of $\chi$ does not play a role, so we allow $\chi$ to be any real number, that is, $\chi\in\mathbb{R}$.
@@
We observe that the well-known “Model H” for incompressible viscous two-phase flows (see [36, 30]) can be recovered from the system (1.1) by neglecting $\sigma$ and setting $\alpha=\beta=\chi=0$. According to (1.7) and (1.9), the resulting Navier–Stokes–Cahn–Hilliard (NSCH) system possesses a dissipative structure (i.e., its total energy decreases over time) and conserves $\overline{\varphi}$ for all $t\geq 0$. Due to these two crucial properties, the NSCH system (with either a regular or a singular potential) has been extensively studied in the literature. In this regard, we refer to [1, 9, 18, 28, 43, 54] and the references therein for the results on well-posedness, long-time behavior of global solutions, and asymptotic analyses. See also [4, 3, 24, 25] for contributions on the more complex case with unmatched densities. When the equation for $\sigma$ is neglected, our system (1.1) also reduces to a variant of the Navier–Stokes–Cahn–Hilliard–Oono (NSCHO) system, which has been studied in [8, 46, 37]. Compared to the basic energy law (1.7), the NSCHO system in [8, 46] exhibits a higher order energy production rate without a definite sign, given by $\alpha\int_{\Omega}(\varphi-c_{0})\widetilde{\mu}\,\mathrm{d}x$ with $\widetilde{\mu}=-\Delta\varphi+\varPsi^{\prime}(\varphi)$ (recalling that $\overline{\varphi}$ depends solely on time and decays to $c_{0}$ exponentially fast). This fact leads to significant challenges in studying global well-posedness and long-time behavior, as discussed in [46]. For the single CHO equation (1.11) without fluid interaction, the authors of [26] overcame the difficulty by considering the Ohta–Kawasaki functional $\mathcal{F}_{\mathrm{OK}}$ and rewriting the CHO equation as
@@
Thanks to the energy balance (1.7) and the mass relations (1.9), (1.10), in our recent work [34], we were able to show that, in the three-dimensional setting, if the initial velocity is small and both the initial phase-field function and the initial chemical density are small perturbations of a local minimizer of the free energy $\mathcal{F}$, then problem (1.1)–(1.3) admits a unique global strong solution. Furthermore, we proved the uniqueness of the asymptotic limit for every global strong solution as $t\to+\infty$ and provided an estimate on the convergence rate. The proofs relied on the dissipative structure of the system and an extended Łojasiewicz–Simon type gradient inequality. However, whether global weak solutions to problem (1.1)–(1.3) exist and will regularize like solutions to the NSCH system, as demonstrated in [1, 3], remains an open question.
@@
(1) Our first result states that every global weak solution $(\bm{v},\varphi,\sigma)$ exhibits a propagation of regularity over time, see Theorem 2.1. Specifically, after an arbitrary positive time, the phase-field $\varphi$ transitions into a strong solution, whereas the chemical density $\sigma$ only partially regularizes. Subsequently, the velocity field $\bm{v}$ becomes regular after a sufficiently large time, followed by a further regularization of the chemical density $\sigma$, which in turn enhances the spatial regularity of $\varphi$. The proof of Theorem 2.1 is inspired by recent work [3] on the global regularity of weak solutions to the NSCH system with unmatched densities. Although we tackle the simpler case of a fluid mixture with constant density, additional challenges arise due to chemotaxis, mass transport, and nonlocal interactions, making it rather difficult to establish the regularity of $(\varphi,\sigma)$. Specifically, the chemical density $\sigma$ exhibits a novel phenomenon, undergoing regularization at different stages of time. For discussions on the strategy of proofs, we refer to Section 2.3.
@@
(2) Our second result pertains to the long-time behavior. Under the additional assumption that $\varPsi$ is real analytic in (1,1)(-1,1) (which is true for (1.5)), we demonstrate that every global weak solution $(\bm{v},\varphi,\sigma)$ converges to a single equilibrium $(\mathbf{0},\varphi_{\infty},\sigma_{\infty})$ as $t\to+\infty$ and provide an estimate of the convergence rate, as stated in Theorem 2.2. This conclusion is reached by applying Theorem 2.1 and employing the same argumentation as presented in [34].
@@
Before concluding the Introduction, we review some related results in the literature. With a regular potential, including (1.6), and some more general reaction terms, problem (1.1)–(1.3) was first studied in [41]. In this work, the authors proved the existence of global weak solutions in both two and three dimensions, and the existence of a unique global strong solution in two dimensions. However, due to the coupling between $\varphi$ and $\sigma$, as well as the loss of maximum principle for the Cahn–Hilliard equation with regular potentials, they imposed certain technical assumptions on the coefficients to achieve their goals. Later, in [31], this restriction was removed by considering a singular potential like (1.5), which is more physically relevant. Using a semi-Galerkin approach, the author proved the existence of global weak solutions in both two and three dimensions, as well as the uniqueness of global weak solutions in two dimensions. For the same system, the authors of [33] further established the global strong well-posedness for arbitrarily large regular initial data in two dimensions. The arguments in [31, 33] rely on the key property that the singular potential $\varPsi$ ensures that the phase function $\varphi$ remains within the physical interval [1,1][-1,1] throughout the time evolution. Moreover, in the two-dimensional case, $\varphi$ is strictly separated from the pure states $\pm 1$ for $t>0$. This separation property plays a crucial role in studying the regularity and long-time behavior of solutions to the Cahn–Hilliard equation with a singular potential, see [5, 12, 26, 45, 47]. We also refer to [1, 9, 14, 24, 27, 28] for generalized systems with fluid coupling. Finally, we note that Oono’s linear term represents possibly the simplest case of a (nonlocal) reaction. Our results provide answers to some open questions raised in [46] concerning the global regularity and asymptotic stabilization of global weak solutions to the NSCHO system with singular potential. On the other hand, more general reaction terms have been integrated into the Cahn–Hilliard equation for biological applications [17, 40], as well as in extended systems for tumor modeling [20, 21, 22, 39]. Investigating the global well-posedness and long-time behavior of global weak/strong solutions to the hydrodynamic system (1.1)–(1.3) with these general reaction terms constitutes an interesting and challenging task.
@@
Plan of the paper. The remainder of this paper is organized as follows. In Section 2, we introduce the mathematical settings and state the main results. In Section 3 we analyze a convective Cahn–Hilliard–diffusion system with a prescribed divergence-free velocity, which may have independent interest. Section 4 is devoted to the proofs of the main results.